OpenStudy (anonymous):

x^2-3x=-10 use the quadratic formula to solve.

OpenStudy (anonymous):

make it equal to zero then substitute the values straight in

OpenStudy (anonymous):

i did that and got a wierd answer,

OpenStudy (anonymous):

it looks like it has complex roots, not real ones http://www.wolframalpha.com/input/?i=x^2-3x%3D-10

OpenStudy (anonymous):

we cant write it like that..

OpenStudy (saifoo.khan):

Molly!!!

OpenStudy (lo):

x^2 - 3x + 10 = 0 then use the quardratic formula to solve. -b^2 plus minus squareroot of b^2 - 4ac , devide by 2a .

OpenStudy (lo):

oh...its been stated >< hee

OpenStudy (anonymous):

i did that but im still confused by it.

OpenStudy (lo):

oh u'd normally get a fraction. or decimal. confusied by wat?

OpenStudy (anonymous):

just th answer and how you get it

OpenStudy (saifoo.khan):

chk your email molly

OpenStudy (lo):

oh. that is the formula. standard formula, if you cant factorize the usual way. the answer you will get from the formula is x itself. got it? =)

jhonyy9 (jhonyy9):

x2 -3x+10=0 x_1,_2=(3+/-sqrt(9-40))/2=(3+/- isqrt(31))/2